OpenStudy (anonymous):

which of the following are vertical asymptotes of the function: y=3cot(2 x)-4 ??

OpenStudy (anonymous):

"of the following"? There are choices? If so, list them.

OpenStudy (anonymous):

1. x=2pi 2. x=pi/3 3. x=+-pi/2 4. x=pi

OpenStudy (paxpolaris):

\[\cot=\frac \cos \sin\]... so, wherever \(\sin(2x)=0 \)

OpenStudy (paxpolaris):

1,3,4

OpenStudy (anonymous):

thank you! :)